I don't know what i got wrong in these. Please help, thanks!

Suppose that x2y3 - cosy - 1. Find dy/dx. dy/dx- Enviar Respuesta Incorrecto. Tries 2/5 Intentos Anteriores This discussion is closed

------

System of Equations Puntos:1 The x-coordinates of two objects moving along the x-axis are given below as function of time t. x1 and x2 never have the same value. Calculate the value of x2 when the objects are nearest to each other. X1 = 21.0t x,--28.0 + 47.0t-10.02 2723 Plot or sketch the two positions vs t. It really helps to understand. Find an expression for the distance they are apart as function of time, then use calculus to get the time when that distance is a minimum. Once you know that t, you can get the positions of the objects Enviar Respueste Incorrecto. Tries 2/5 Intentos Anteriores

------

Vector operations Puntos:1 The labelled vectors below are drawn to scale For each of the statements select Greater than, Less than, or Equal to The magnitude of S... that of H H.Y is... 0 Is x ZI is 0. Y N is 0. IR × SI is 0. Enviar Respuesta Tries 0/5
Answer #1

1.

2.3 t y - cos y-1

2xy^3+3x^2y^2rac {dy}{dx}+sin yrac {dy}{dx}=0

2xy 32.2 y2 + sin y di

2.

r(t)-x1-T2 = 10t2-261+ 28

dr(t) dt =20t-26

dr(t) dt 20t-26 = 0

Rightarrow t=1.3: s

fr(t) = 20 > 0 dt2

T2(t = 1.3 s) = 16.2

3.

(i) The magnitude of S is less than H

(ii) Since H is perpendicular to Y, H dot Y is zero.

(iii) It is less than zero.

(iv) Since S and Z are parallel to each other, their cross product is zero.

(v) Greater than zero.

(vi) It is less than zero since the angle between R and Z is obtuse.